OpenStudy (anonymous):

can someone check my work? Given: f(x) = x^3 - x^2 - 20x + 4 Find all numbers "c" that verify Rolles THM on the interval [0, 5]. My work will be in comments.

OpenStudy (anonymous):

f'(x) = 3x^2 - 2x - 20 set that = 0; 3x^2 - 2x - 20 = 0 .................x(3x-2) = 20

OpenStudy (anonymous):

I get x = 0 and x = 2/3 ... but apparently that is not right?

OpenStudy (anonymous):

f(x) is continuous and differentiable everywhere and f(0) = f(5)